EXISTENCE AND REGULARITY OF A WEAKLY HARMONIC MAP INTO A FINSLER SPACE WITH A SPECIAL STRUCTURE

ATSUSHI TACHIKAWA

Abstract : In this talk we treat the existence and interior Hölder regularity of an energy minimizing weakly harmonic map from a m-dimensional Riemannian manifold (M, g) into a n-dimensional Finsler space (N, F) which is given as a modification of a Riemannian manifold (N, h).